OpenStudy (anonymous):

Let f(x)=(x+6)/5. Find f^-1(6)

OpenStudy (slaaibak):

y = (x+6)/5 for the inverse: x = (y+6)/5 5x - 6 = y f^-1(16) = 5 * 16 - 6 = 74

OpenStudy (slaaibak):

Oh, f^-1(6).. I misread, sorry. Then it's f^-1(6) = 5 * 6 - 6 = 24

OpenStudy (mertsj):

Find f(6) and then switch the x and y

OpenStudy (anonymous):

okay when i find f(6) i get 12/5... then what?

OpenStudy (mertsj):

that means the ordered pair (6,12/5) is part of the function f so (12/5,6) is a member of f^-1

OpenStudy (slaaibak):

Actually, you can set 6 = (x+6)/5 Then the x = f^-1(6) but rather don't do that, since you have to fully grasp why that is possible

OpenStudy (jamesj):

@mertsj, Even so, that doesn't tell you the value of f^-1(6). A simple way to think about this problem is the value or values of f^-1(6) is: What are the value or values of x for which f(x) = 6? Hence f(x)=(x+6)/5 = 6. Now solve for x.

OpenStudy (slaaibak):

\[f(x) = y \rightarrow x = f^{-1}(y)\]